OpenStudy (anonymous):

Solve. Write the answer in simplest form. (Must show work!) 1. 3/4m = 5/6

OpenStudy (texaschic101):

you can either divide both sides by 3/4 or multiply the equation by the LCD, which is 12 and solve for m. Either way you should get the same answer.

OpenStudy (anonymous):

I need the work shown.

OpenStudy (texaschic101):

which way you want to do it.....dividing by 3/4 or the other way ?

OpenStudy (anonymous):

i guess 3/4 idk

OpenStudy (anonymous):

nd my answer is 1 1/9 idk if its correct

OpenStudy (texaschic101):

3/4m = 5/6 (divide both sides by 3/4) m = (5/6) / (3/4) m = 5/6 * 4/3 m = 20/18 reduces to 10/9 or 3/4m = 5/6 (multiply by 12) 9m = 10 (divide by 9) m = 10/9 you can pick which way you want to show your work

OpenStudy (anonymous):

so 10/9 is the answer ?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

is 1 1/9 wrong ?

OpenStudy (texaschic101):

no...if you want it in mixed number form, 1 1/9 is correct. I believe you could put 10/9 or 1 1/9 and both answers are correct, unless the question specifically wants it a certain way

OpenStudy (anonymous):

Oh can u write on the drawing pad thing cus i cant tell the "/" meant divide or fractions

OpenStudy (texaschic101):

it means fraction. Although, if you divide, it turns it into a decimal.

OpenStudy (anonymous):

So all the "/" means fractions and not either of them is dividing

OpenStudy (texaschic101):

well..... 5/6(fraction) divided by 3/4(fraction)

OpenStudy (anonymous):

Oh i get it now, thanks btw i undestand now

OpenStudy (texaschic101):

good to hear :) glad I could help